OpenStudy (anonymous):

Suppose that |2x-3|≤5. What are the minimum and maximum possible values of |3x-7|?

OpenStudy (anonymous):

How do I go on solving this?

OpenStudy (math&ing001):

Well first we find minimum and maximum possible values of x.

OpenStudy (math&ing001):

Can you do it ?

OpenStudy (math&ing001):

@drag0night

OpenStudy (anonymous):

i can do it for |2x-3|≤5

OpenStudy (anonymous):

x = 4 and x = -1

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

Yeah those are the max and min values of x. - 1 <= x <= 4 -3 <= 3x <= 12 4 <= 3x+7 <= 19

OpenStudy (anonymous):

ohh that makes sense!!!

OpenStudy (anonymous):

Thank you :D

OpenStudy (math&ing001):

Welcome =)

OpenStudy (anonymous):

i don't know why but my textbook gives the right answer as minimum = 0 and max = 10 @math&ing001

OpenStudy (math&ing001):

I don't see how that's possible, maybe a mistake.

OpenStudy (math&ing001):

Oh wait a sec !

OpenStudy (math&ing001):

It's -7 not +7. - 1 <= x <= 4 -3 <= 3x <= 12 -10 <= 3x-7 <= 5 <=10 |3x-7|<=10

OpenStudy (anonymous):

Ok now everything makes sense. thanks for your time!